OpenStudy (anonymous):

im suppose to simplify this 4y-1/2y^2-3y-2 2y-1/y-2 with the restirctions y cannot equal -1/2 and y cannot equal -2...is that right?

OpenStudy (anonymous):

you want to simplify this: \[\dfrac{4y-1}{2y^2-3y-2} \] ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh but its 4^2

OpenStudy (amistre64):

as is, the top is a clue to how youd want to try to factor the bottom

OpenStudy (amistre64):

.... then that makes the top a difference of squares that way :)

OpenStudy (amistre64):

still a good clue tho as to how you should try to factor the bottom tho

OpenStudy (anonymous):

could you please retype the task? i still don't understand what exactly should be simplifyed=(

OpenStudy (anonymous):

u wrote it right its just 4y squared. the whole thing needs to be simplfied

OpenStudy (amistre64):

\[\dfrac{4y^2-1}{2y^2-3y-2}\]

OpenStudy (anonymous):

\[\dfrac{4y^2-1}{2y^2-3y-2}=\dfrac{(2y-1)(2y+1)}{(2y+1)(y-2)}=\dfrac{2y-1}{y-2} y \neq-1/2\]

OpenStudy (anonymous):

So you were right=)

OpenStudy (anonymous):

ohh yay! thanks !